OpenStudy (anonymous):

Jason places his economics book (3 kg) on a level table. He then puts his calculus book (5 kg) on top of the economics book, then puts his physics book (2 kg) on top of the calculus book. Find the force that the table applies to the economic book.

OpenStudy (anonymous):

All the bodies are on equilibrium. Due to Newton's third law, the table applies the same amount of force from the 3 books, but on the opposite direction, on the fist book (since the table only interacts directly with it). That force is equal to the mass of all the books multiplied by the acceleration of gravity: 10*g

OpenStudy (shamim):

I eanna explain it more!!!

OpenStudy (shamim):

Total mass of books=3+5+2=?

OpenStudy (shamim):

Weight of thisc10kg book is=mg=10*9.8=98N Right?

OpenStudy (shamim):

Actually this force is applied on table by books

OpenStudy (shamim):

And the table is also exerted the force 98N on the economics book

OpenStudy (shamim):

Because every action has an equal n oposit reaction